# Game Theory in Cosmology

### v1

@@ Abstract
We present a game-theoretic statistical framework for cosmology, which we term Cosmological Teleodynamics. We recast the dark sector, cosmic acceleration, large-scale structure, and cosmic tensions as emergent consequences of nonlocal memory and intrinsically persistent organization in a self-gravitating Universe. By introducing a maximum-caliber weight on cosmic histories and a bias functional encoding structural memory, we derive modified Friedmann, Boltzmann, and Poisson equations that naturally generate dark energy-like acceleration, dark matter-like clustering, and scale-dependent growth suppression. We also show how this approach can naturally help alleviate the $H_{0}$ and $S_{8}$ tensions, can produce anisotropic velocity fields, and predict environment-dependent halo signatures that cannot arise from particle dark matter or scalar-field dark energy. We also derive a generalized horizon entropy and temperature, revealing a nonequilibrium statistical origin for cosmic acceleration and formulating a Law of Universal Arbitrage Equilibrium that governs the evolution of the Universe, showing that it is expanding towards a continuous form of Nash equilibrium. Cosmological Teleodynamics therefore offers a unified, emergent, and testable alternative to the conventional dark sector, rooted not in new particles but in the intrinsic statistical and systemic structure of cosmic memory, and shows that the universe itself operates like a giant potential game.
@@ 1 Introduction
To address these concerns, we present a perspective that accounts for persistence effects in a novel manner by using game theory. One might wonder how game theory applies to cosmology. It may seem rather far-fetched. However, as we demonstrate, there is a surprisingly close connection via the concept of persistence. Game theory is concerned with predicting the behavior of goal-driven agents that, by their very nature of incessantly pursuing goals, exhibit persistent behavior. This persistence results in constrained behavior - i.e., persistent agents do not explore the entire phase space but only a part of it. Thus, persistence implies latent constraints on the dynamics and evolution of agents. Game theory offers us a new way to model persistence mathematically as constraints on possible collective behaviors. It is in this sense that we use the word ‘game’ in our theory of cosmology.
@@
Thus, we present a teleodynamic model of the universe whose evolution is not purely random but is influenced by persistent structures, memory effects, and dynamically induced preferences. In conventional statistical mechanics, we assume that the probability assigned to a microscopic configuration is determined solely by its energy via the Boltzmann weight. We further assume ergodicity, loss of memory, short-range interactions, etc. These assumptions are questionable for gravitational and cosmological systems where correlations are long-ranged, ergodicity is broken, and macroscopic evolution depends sensitively on prior structure formation. In this sense, teleodynamics expands the statistical toolkit by adding a second structural ingredient, in addition to energy, namely a functional encoding of the system’s persistent organizational bias.
@@ 2 Potential Games and Arbitrage Equilibrium
${h}_{j}(\boldsymbol{x})\equiv{\partial\phi(\boldsymbol{x})}/{\partial x_{j}}$ (1)
@@
One can view the thermodynamics of gas molecules in a container via the lens of potential games - the Thermodynamic Game [40]- and define the effective “utility” $h_{j}$ of a molecule in state $j$ as
$h_{j}=-\beta E_{j}-\ln N_{j}$ (5)
where $E_{j}$ is its energy, $N_{j}$ is the number of molecules in state $j$, and $\beta=1/k_{B}T$ is the Boltzmann factor. Using Eq. 5 in Eq. 2, we have
$\phi(\textbf{x})=-\beta\sum_{j=1}^{n}x_{j}E_{j}+\frac{1}{N}\ln\frac{N!}{\prod_{j=1}^{n}(N_{j})!}$ (6)
Using Eq. 6 in Eq. 3, we obtain the well-known Gibbs-Boltzmann exponential probability distribution at equilibrium as
$x_{j}^{*}=\frac{e^{-\beta E_{j}}}{\sum_{j=1}^{n}e^{-\beta E_{j}}}$ (7)
@@
$x_{j}^{*}=\frac{1}{Z}e^{-\beta E_{j}}$ (9)
Readers familiar with statistical mechanics would immediately recognize that the second term in Eq. 6 corresponds to entropy (SS[35]. Therefore, as observed by Venkatasubramanian [35, 41], we recognize that by maximizing $\phi$, we are indeed maximizing the entropy $S$ subject to the energy constraint given by the first term in Eq. 6.
@@
$\tilde{\phi}(\boldsymbol{x})=\sum_{j=1}^{n}\int\left[u_{j}-v_{j}-\ln x_{j}\right]dx_{j}$ (11)
@@
$x_{j}^{*}=\frac{1}{\tilde{Z}}\exp(u_{j}-v_{j})$ (13)
As shown in Eq. 6, the $\ln N_{j}$ term leads to entropy $S$ upon integration, and the $u_{j}$ and $v_{j}$ terms (upon integration) correspond to the constraints that maximum entropy must satisfy. Thus, the persistence preferences of goal-driven agents, encoded in the terms $u_{j}$ and $v_{j}$, manifest themselves as constraints on maximum entropy. From Table 1, we have (i) for gas molecules, $u_{j}=0$ and $v_{j}=\beta E_{j}$, (ii) for income distribution in economics, $u_{j}=\alpha\ln S_{j}$ and $v_{j}=\eta(\ln S_{j})^{2}$, and (iii) for segregation dynamics in sociology, $u_{j}=\alpha N_{j}+\ln(H-N_{j})$, $v_{j}=\eta N_{j}^{2}$. These different mathematical forms of constraints (i.e., $\sum_{j=1}^{n}\int\left[u_{j}-v_{j}\right]dx_{j}$) for maximum entropy across various domains show how persistence effects can appear mathematically in our framework. Eq. 12 and 13 could get complicated if $u_{j}$ and $v_{j}$ depend on $N_{j}$, as in the case of bird flocking and social segregation games shown in Table 1. However, this is not the case for the Thermodynamic and Income Games. As a result, we can obtain the probability distributions analytically, cleanly - e.g., exponential (energy) and lognormal (income) distributions.
@@ 3 Statistical Teleodynamics of Cosmology
Guided by the mathematical models of persistence in other domains, as shown in Table 1, we are now ready to develop a teleodynamic framework for cosmology that captures the effects of persistent structures, memory effects, and dynamically induced preferences. We start with Eq. 10, Eq. 12, and Eq. 13, considering a discrete space of microstates $x$. Instead of assigning weights solely by the energy $E(x)$, as in Eq. 5, we introduce an additional functional $\Phi(x)$ which captures a path-dependent or environment-dependent persistence bias as a constraint intrinsic to the system. Therefore, we have $u(x)=0$ and $v(x)=\beta E(x)+\alpha\Phi(x)$, yielding
$h(x)=-\beta E(x)-\alpha\Phi(x)-\ln N(x)$ (14)
@@
$p(x)=\frac{1}{Z}\exp[-\beta E(x)-\alpha\Phi(x)]$ (15)
@@
$Z(\beta,\alpha)=\sum_{x}\exp[-\beta E(x)-\alpha\Phi(x)]$ (16)
@@
Since it is natural to use free energy instead of the game-theoretic (i.e., teleodynamic) potential in cosmology, we will now switch to a free energy-based analysis. As Venkatasubramanian showed [35, 34, 41], minimizing free energy is equivalent to maximizing game-theoretic potential $\phi$.
So, when $\Phi$ is present, the effective free energy becomes
$F_{\text{TD}}=-\frac{1}{\beta}\ln Z=\sum_{x}p(x)E(x)+\frac{1}{\beta}\sum_{x}p(x)\ln p(x)+\alpha\sum_{x}p(x)\Phi(x)$ (17)
This tells us that the equilibrium distribution simultaneously minimizes an energetic, an entropic, and a teleodynamic term. The third contribution has no analog in ordinary thermodynamics and encodes the persistence of organizational constraints and dynamical feedback not reducible to energy alone.
@@
$\dot{x}=-\nabla E(x)-\alpha\nabla\Phi(x)+\eta(t)$ (18)
@@
$\frac{\partial p(x,t)}{\partial t}=\nabla\cdot\left[p(x,t)\nabla\left(E(x)+\alpha\Phi(x)\right)\right]+D\nabla^{2}p(x,t)$ (19)
@@
The structure of this whole theory consists of a few vital ingredients: first is a generalized Gibbs distribution that retains the form of an exponential family but includes an additional functional in the exponent. The second is the appearance of an effective potential $E_{\text{eff}}(x)=E(x)+\alpha\Phi(x)$ that modifies the free energy landscape and the deterministic drift of coarse-grained dynamics. The third is the recognition that complex macroscopic systems may not be described by equilibrium thermodynamics, but rather by a combination of free-energy minimization and attractor selection, governed by $\Phi$. These features become particularly important in cosmology, where long-range gravitational forces, nonlocality, and the presence of large-scale correlated structures imply that the assumptions of ordinary statistical mechanics are not satisfied. In such a context, the teleodynamic corrections provide a natural mechanism for introducing effective energy densities, biasing large-scale flows, and generating attractor behavior without invoking new particle species or additional fundamental fields. The key differences between thermodynamics and teleodynamics are summarized in Table 2.

### v2

@@ Abstract
We present a game-theoretic statistical framework for cosmology, which we term Cosmic Teleodynamics. We recast the dark sector, cosmic acceleration, large-scale structure, and cosmic tensions as emergent consequences of nonlocal memory and intrinsically persistent organization in a self-gravitating universe. Our main insight is that, just as the Big Bang and Inflation left their signature on the Cosmic Microwave Background (CMB), they also left a similar lasting imprint on the spacetime itself, influencing future dynamics and evolution. By introducing a maximum-caliber weight on cosmic histories and a bias functional encoding structural memory, we derive modified Friedmann, Boltzmann, and Poisson equations that naturally generate dark energy-like acceleration, dark matter-like clustering, and scale-dependent growth suppression. We also show how this approach can naturally help alleviate the $H_{0}$ and $S_{8}$ tensions, can produce anisotropic velocity fields, and predict environment-dependent halo signatures that cannot arise from particle dark matter or scalar-field dark energy. We also derive a generalized horizon entropy and temperature, revealing a nonequilibrium statistical origin for cosmic acceleration. We formulate a Law of Universal Arbitrage Equilibrium that governs the evolution of the Universe, showing that it is expanding towards a continuous form of Nash equilibrium. We also show that our approach naturally explains the Coincidence problem as well. Cosmic Teleodynamics therefore offers a unified, emergent, and testable alternative to the conventional dark sector, rooted not in new particles but in the intrinsic statistical and systemic structure of cosmic memory. The crux of this framework is, by just considering a more complete form of statistical mechanics, one can address multiple major problems in modern cosmology.
@@ 1 Introduction
To address these concerns, we present a perspective that accounts for persistence effects in a novel manner by using game theory. One might wonder how game theory applies to cosmology. It may seem rather far-fetched. However, as we demonstrate, there is a surprisingly close connection via the concept of persistence. Game theory is concerned with predicting the behavior of goal-driven agents that, by their very nature of incessantly pursuing goals, exhibit persistent behavior. This persistence results in constrained behavior, where persistent agents do not explore the entire phase space but only a subset of it. Thus, persistence implies latent constraints on the dynamics and evolution of agents. Game theory provides a new approach to modeling persistence mathematically as constraints on possible collective behaviors. It is in this sense that we use the word ‘game’ in our theory of cosmology.
@@
Thus, we present a teleodynamic model of the universe, which we call Cosmic Teleodynamics, whose evolution is not purely random, but is influenced by persistent structures, memory effects, and dynamically induced preferences. In conventional statistical mechanics, one assumes that the probability assigned to a microscopic configuration is determined solely by its energy via the Boltzmann weight. One further assumes ergodicity, loss of memory, short-range interactions, etc. These assumptions are questionable for gravitational and cosmological systems where correlations are long-ranged, ergodicity is broken, and macroscopic evolution depends sensitively on prior structure formation. In this sense, teleodynamics expands the statistical toolkit by adding a second structural ingredient, in addition to energy, namely a functional encoding of the system’s persistent organizational bias.
@@ 2 Potential Games and Arbitrage Equilibrium
${h}_{j}(\boldsymbol{x})\equiv{\partial\phi(\boldsymbol{x})}/{\partial x_{j}}$ (2.1)
@@
One can view the thermodynamics of gas molecules in a container via the lens of potential games - the Thermodynamic Game [40]- and define the effective “utility" $h_{j}$ of a molecule in state $j$ as
$h_{j}=-\beta E_{j}-\ln N_{j}$ (2.5)
where $E_{j}$ is its energy, $N_{j}$ is the number of molecules in state $j$, and $\beta=1/k_{B}T$ is the Boltzmann factor. Using Eq. 2.5 in Eq. 2.2, we have
$\phi(\textbf{x})=-\beta\sum_{j=1}^{n}x_{j}E_{j}+\frac{1}{N}\ln\frac{N!}{\prod_{j=1}^{n}(N_{j})!}$ (2.6)
Using Eq. 2.6 in Eq. 2.3, we obtain the well-known Gibbs-Boltzmann exponential probability distribution at equilibrium as
$x_{j}^{*}=\frac{e^{-\beta E_{j}}}{\sum_{j=1}^{n}e^{-\beta E_{j}}}$ (2.7)
@@
$x_{j}^{*}=\frac{1}{Z}e^{-\beta E_{j}}$ (2.9)
Readers familiar with statistical mechanics would immediately recognize that the second term in Eq. 2.6 corresponds to entropy (SS[35]. Therefore, as observed by Venkatasubramanian [35, 41], we recognize that by maximizing $\phi$, we are indeed maximizing the entropy $S$ subject to the energy constraint given by the first term in Eq. 2.6.
@@
$\tilde{\phi}(\boldsymbol{x})=\sum_{j=1}^{n}\int\left[u_{j}-v_{j}-\ln x_{j}\right]dx_{j}$ (2.11)
@@
$x_{j}^{*}=\frac{1}{\tilde{Z}}\exp(u_{j}-v_{j})$ (2.13)
As shown in Eq. 2.6, the $\ln N_{j}$ term leads to entropy $S$ upon integration, and the $u_{j}$ and $v_{j}$ terms (upon integration) correspond to the constraints that maximum entropy must satisfy. Thus, the persistence preferences of goal-driven agents, encoded in the terms $u_{j}$ and $v_{j}$, manifest themselves as constraints on maximum entropy. From Table 1, for example, we have (i) for gas molecules, $u_{j}=0$ and $v_{j}=\beta E_{j}$, (ii) for income distribution in economics, $u_{j}=\alpha\ln S_{j}$ and $v_{j}=\eta(\ln S_{j})^{2}$, and (iii) for segregation dynamics in sociology, $u_{j}=\alpha N_{j}+\ln(H-N_{j})$, $v_{j}=\eta N_{j}^{2}$. These different mathematical forms of constraints (i.e., $\sum_{j=1}^{n}\int\left[u_{j}-v_{j}\right]dx_{j}$) for maximum entropy across various domains show how different persistence effects can appear mathematically in our framework. Eq. 2.12 and 2.13 could get complicated if $u_{j}$ and $v_{j}$ depend on $N_{j}$, as in the case of bird flocking and social segregation games shown in Table 1. However, this is not the case for the Thermodynamic and Income Games. As a result, we can obtain the probability distributions analytically, cleanly - e.g., exponential (energy) and lognormal (income) distributions.
@@
3 Persistence and Memory in Cosmology
Building on our statistical teleodynamic formulation of persistence and memory, in this section, we define them in a cosmological context. Although the Big Bang is not like an explosion, the following example is nonetheless intuitively helpful. Consider a familiar, if disturbing, account of a bomb going off in the middle of a large crowd of people. As soon as the bomb explodes, people start scattering in different directions. Several minutes after the explosion, people are still trying to escape in whatever way they can. Some are running, some are jumping into a car, some are riding a bike, some are trying to hide in a building, and so on. To remain safe, people do different things depending on what their local constraints allow. If the destruction is massive, people may try to escape to safety even hours or days after the explosion. Thus, days, months, or even years after a catastrophic event, people remember the event and its consequences. The event creates a generic systemic memory that persists over time. It also creates specific memories for different people of how they reacted, how they escaped, and how it forever changed their lives. They all share the same systemic memory of the explosion, but they all remember the specific details differently based on their individual experiences. The point of this example is that catastrophic events leave both generic and specific signatures on the elements of a system that persist for a long time.
Similarly, we suggest that the Big Bang and subsequent inflationary events (BBI) are of such cosmic magnitude, both literally and figuratively, that they left a persistent memory in the dynamics and evolution of the universe. The sudden inflationary expansion by about $10^{30}$ in scale, which lasted only from $10^{-36}$ to $10^{-32}$ seconds, was such an extraordinary cosmic shock that the universe still "remembers" it. The intrinsic quantum fluctuations that occurred during BBI produced an uneven spacetime structure. These non-uniformities persisted and became enormously magnified during inflation and the subsequent expansion of the universe. These later evolved into galactic proportions seen in the current macroscopic structure of the universe. Our main insight is that the BBI not only left its signature in the Cosmic Microwave Background (CMB), but also left a similar, indelible, uneven structures on the very fabric of the cosmos, namely, the spacetime.
We propose that the BBI left two kinds of persistent structural signatures on the spacetime fabric, one generic and one specific. Every element of spacetime remembers the BBI, but each in its own way. The generic signature is like a background effect that influences everything in the same manner, while the specific signature allows stellar objects to evolve according to their local histories and constraints. This is what we model as the teleodynamic bias potential $\Phi(t,x)$ in the next section, which we decompose into its generic background part ($\bar{\Phi}(t)$) and the specific fluctuations part ($\varphi(t,x)$), given by
$\Phi(t,x)=\bar{\Phi}(t)+\varphi(t,x)$ (3.1)
In the teleodynamic framework, memory ($\Phi(t,x)$) is not a psychological or informational construct, but a physical property of an evolving cosmic system. It represents the persistent imprint left on the large-scale gravitational field by the BBI, the history of structure formation, matter flows, and tidal interactions across cosmic timescales. Because gravity in an expanding universe is long-range and non-ergodic, the distribution of matter does not “forget” its past configurations, including the BBI, which is to say that we cannot treat the physics here to be completely Markovian. Instead, every merger, filamentary collapse, void expansion, or tidal alignment leaves traces in the statistical and systemic organization of the cosmic web, and these persistent features accumulate into a global bias functional, which we refer to as cosmic memory.
From a dynamical perspective, cosmic memory encodes how the universe departs from the assumptions of perfect equilibrium that underlie standard statistical mechanics. This is exemplified by the fact that a collisionless, self-gravitating system never fully thermalizes, and its phase-space distribution maintains long-lived correlations that cannot be erased by expansion alone. These correlations are precisely what teleodynamics captures. As the universe evolves, preferred organizational patterns survive Hubble timescales and feed back into both the background expansion and the growth of structure. In this sense, we say that memory arises as an emergent, non-local statistical field that quantifies how past gravitational organization biases affect evolution.
Such process-dependent memory effects on macroscopic structure and properties are not unique to cosmology. The formation and evolution of macroscopic structures in materials science have similar features. For example, in the design, formulation, and manufacturing of rubber-based engineered materials, small microscopic irregularities, such as fluctuations in temperature and stress profiles during the curing process, can transform into large macroscopic variations in structure, properties, and behavior. This micro-to-macro mathematical modeling is complicated by interactions between physical and chemical transformations arising from complex sulfur-chemistry reactions, as discussed by Ghosh et al. [42]. Of course, one needs to be careful, as all analogies have limitations. However, structural formation and evolution in material science can provide valuable insight into what to expect in other domains.
Cosmologically speaking, this memory becomes especially important once structure formation becomes nonlinear, as is the case in the current age of the universe. Filaments, clusters, and voids act as reservoirs of persistent tidal information, storing the integrated effects of billions of years of gravitational interactions. This accumulated organization cannot be described by local density alone; instead, it manifests itself as a global functional of the cosmic web. Viewed this way, cosmic memory is not an additional field or exotic substance introduced by hand, but is instead the intrinsic statistical and systemic outcome of a universe that has evolved under long-range gravity, formed a hierarchy of structures, and retained correlations across cosmic timescales. Teleodynamics formalizes this insight into a predictive framework: the accumulated memory of structure formation influences large-scale dynamics, shapes late-time acceleration, and modulates the growth of perturbations. In this sense, cosmic memory is the physical bridge connecting the microscopic history of gravitational organization and the macroscopic observables that define modern cosmology.
@@ 4 Statistical Teleodynamics of Cosmology
Guided by these considerations and the mathematical models of persistence and memory in other domains, as shown in Table 1, we are now ready to develop a teleodynamic framework for cosmology that captures the effects of persistent structures, memory effects, and dynamically induced preferences. We start with Eq. 2.10, Eq. 2.12, and Eq. 2.13, considering a discrete space of microstates $x$. Instead of assigning weights solely by the energy $E(x)$, as in Eq. 2.5, we introduce an additional functional $\Phi(x)$ which captures a path-dependent or environment-dependent persistence bias as a constraint intrinsic to the system. Therefore, we have $u(x)=0$ and $v(x)=\beta E(x)+\alpha\Phi(x)$, yielding
$h(x)=-\beta E(x)-\alpha\Phi(x)-\ln N(x)$ (4.1)
@@
$p(x)=\frac{1}{Z}\exp[-\beta E(x)-\alpha\Phi(x)]$ (4.2)
@@
$Z(\beta,\alpha)=\sum_{x}\exp[-\beta E(x)-\alpha\Phi(x)]$ (4.3)
@@
Since it is natural to use free energy instead of the game-theoretic potential in cosmology, we will now switch to a free energy-based analysis. As Venkatasubramanian showed [35, 34, 41], minimizing free energy is equivalent to maximizing game-theoretic potential $\phi$.
To make the structure of the teleodynamic free energy ($F_{\rm TD}$) explicit, it is useful to derive it directly from the microscopic probability distribution. We begin with the expression (4.2) which implies
$\ln p(x)=-\ln Z-\beta E(x)-\alpha\Phi(x)$ (4.4)
Multiplying this expression by $p(x)$ and summing over all microstates gives
$\sum_{x}p(x)\ln p(x)=\sum_{x}p(x)\big[-\ln Z-\beta E(x)-\alpha\Phi(x)\big]$ (4.5)
Using the normalization condition $\sum_{x}p(x)=1$, this becomes
$\sum_{x}p(x)\ln p(x)=-\ln Z-\beta\sum_{x}p(x)E(x)-\alpha\sum_{x}p(x)\Phi(x)$ (4.6)
Solving this relation for $\ln Z$ yields
$\ln Z=-\sum_{x}p(x)\ln p(x)-\beta\sum_{x}p(x)E(x)-\alpha\sum_{x}p(x)\Phi(x)$ (4.7)
Inserting this into the definition of the teleodynamic free energy $F_{\rm TD}=-\frac{1}{\beta}\ln Z$, gives us the following
$F_{\rm TD}=\frac{1}{\beta}\sum_{x}p(x)\ln p(x)+\sum_{x}p(x)E(x)+\frac{\alpha}{\beta}\sum_{x}p(x)\Phi(x)$ (4.8)
This expression makes clear that the equilibrium distribution minimizes a combination of energetic, entropic and teleodynamic contributions, with the final term encoding the effects of persistent memory and structural bias that have no analogue in ordinary thermodynamics. The third contribution has no analog in ordinary thermodynamics and encodes the persistence of organizational constraints and dynamical feedback not reducible to energy alone.
@@
$\dot{x}=-\nabla E(x)-\alpha\nabla\Phi(x)+\eta(t)$ (4.9)
@@
$\frac{\partial p(x,t)}{\partial t}=\nabla\cdot\left[p(x,t)\nabla\left(E(x)+\alpha\Phi(x)\right)\right]+D\nabla^{2}p(x,t)$ (4.10)
@@
The structure of this whole theory consists of a few vital ingredients: first is a generalized Gibbs distribution that retains the form of an exponential family but includes an additional functional in the exponent. The second is the appearance of an effective potential $E_{\text{eff}}(x)=E(x)+\alpha\Phi(x)$ that modifies the free energy landscape and the deterministic drift of coarse-grained dynamics. The third is the recognition that complex and evolving macroscopic systems may not be described by equilibrium thermodynamics, but rather by a combination of free-energy minimization and attractor selection, governed by $\Phi$. These features become particularly important in cosmology, where long-range gravitational forces, nonlocality, and the presence of large-scale correlated structures imply that the assumptions of ordinary statistical mechanics are not satisfied. In such a context, the teleodynamic corrections provide a natural mechanism for introducing effective energy densities, biasing large-scale flows, and generating attractor behavior without invoking new particle species or additional fundamental fields. The key differences between thermodynamics and teleodynamics are summarized in Table 2 and in Eq. 2.11-2.13.